solve the system of linear eqations by sybstitution. x+2y=6 and 2x+3y=8
i do have the results which are -2,4. But how do i find out type of system, and type of lines,

Answer by ewatrrr(24785)

 
Hi,
x+2y=6 Green
2x+3y=8 Blue
-2x-4y=-12
2x+3y=8
-y = -4
y = 4 and x = -2
|Yes! Ordered pair Pt(-2,4)is the solution for this INDEPENDENT system:See below
Lines intersect at Pt(-2,-4)

A consistent system has at least one solution.
An inconsistent system has no solutions.
An independent system has exactly one solution. ***
A dependent system has infinitely many solutions.